Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 miles per hour and train B is traveling at 48 miles per hour. Train A passes a station at 3:10 A.M. If train B passes the same station at 3:25 A.M, at what time will train B catch up to train A??

When will train B catch up to train A??

Answer by checkley77(12844)

d=rt
Train a: d=40t
Train b: d=48(t-.25)
40t=48(t-.25)
40t=48t-12
40t-48t=-12
-8=-12
t=-12/-8
t=1.5 hours after 3:10 they will meet.
Proof:
40*1.5=48(1.5-.25)
60=48*1.25
60=60